A charge of q1=3.6mC is located at (x=288, y=-182)cm. A charge of q2=4.3mC is located at (3.0,3.8)m. A charge of q3=9.6mC is located at (-34.2,-13)m. What is the y-component of the force on q1 due to q2and q3? Thanks for helping!
Answer #1

A charge of q1-3.6mcis located at x=288, y=-182 cm. A charge of q2-4.3mcis located at due to q2 and q3? 3.0 3.8 m. A charge o

3-2.83)+.82-(-1,82 (56 ч)2 2 (2, 88-(34,2))2十(-1,82-(-13) )ユ 38.73 13 2.88-(-34,2) 1.813) tan OB= 13 2. 2. 38.1

angle iwth y-axis 1 2- an 2.883 Op =1,23 2. Net force in Y-direction